# The Machine Learning Control Method for Counterfactual $*{}^{*}$

### v1

@@ 4.3.1 Design Stage
Table 3 below reports the average performance—in terms of MSE—of the four selected ML algorithms in forecasting the standardized math test score of fifth-grade students across the pre-treatment test sets (2018 and 2019).
The best-performing MLCM version is the one using the random forest with 1/4 of the number of variables as candidates at each split. Overall, the fully non-linear random forest and boosting fare better than the linear ML methods, suggesting the presence of non-linearities in the data-generating process. Moreover, the non-linear methods outperform LASSO and Partial Least Squares in all cases, and their performance is stronger in the presence of very few variables and less deteriorated by the inclusion of variables with poor predictive power (cf. Table 12 in Appendix E). The optimal number of variables for random forest corresponds to the 20 variables with the highest importance score attributed by the preliminary forest (see Table 13 in Appendix E for a detailed description of these variables).
Importantly, panel CV on pre-treatment period also implicitly allows one to carry out an in-time placebo analysis along the lines of Bertrand et al., (2004) and Liu et al., (2022). In-time placebos are performed on the same pool of treated units where we “fake” that the treatment occurred at time $t_{0}-1$ and use only information up to $t_{0}-1$ to forecast the counterfactuals at time $t_{0}$. As we know what the real values at time $t_{0}$ are, we can use the difference between the actual values and the forecasted values to evaluate forecasting accuracy and assess the unbiasedness of our estimator, as well as the distribution and spatial autocorrelation of the residuals. Put simply, we shift the treatment date artificially backward in time and assess the difference with the main estimates.
As reported in the table, the placebo ATEs are indistinguishable from zero and statistically insignificant both in 2018 (ATE: 0.0554, with 95% confidence intervals (–0.1182; 0.2765)) and 2019 (ATE: 0.0254, with 95% confidence intervals (–0.0273; 0.0728)). [^25] The placebo effects remain statistically indistinguishable from zero when using default hyperparameter values. By “default”, we refer to hyperparameters set at central values based on the most predictive variables—specifically, 20 variables with the highest importance scores from the preliminary forest, and 1/3 of the total number of variables as candidates at each split. Similarly, the counterfactual estimates exhibit substantial similarity regardless of whether a preliminary random forest or an alternative preliminary procedure is employed, such as boosting or selecting covariates based solely on the highest absolute correlation coefficients (e.g., the correlation between the counterfactual estimates and those obtained using default hyperparameter values is 0.9985). Detailed results are available upon request.
@@
Panel A - Performance
MLCM using: MSE
@@
Panel B – Placebo estimates
@@
Notes. The panel CV procedure has selected the 20 most predictive variables out of the initial dataset for all methods. In addition, for boosting, it has selected 2,000 trees, 2 as the maximum depth of each tree, 5 as the minimum number of observations in terminal nodes and 0.005 for the learning rate. The other hyperparameters selected are: 0.3 as the penalty parameter (LASSO), 9 as the number of components (PLS), and 6 for the number of variables randomly sampled as candidates at each split (random forest). Lower and upper bounds of placebo ATE estimates refer to the 95% confidence interval and are reported within brackets.
In addition, the unit-level placebo map in Figure 2 depicts the temporal average (for the years 2018 and 2019) individual ‘treatment’ effects estimated with the best-performing algorithm, random forest, and shows that almost all LLMs exhibit no trace of significant differences between the forecasted and observed math scores in the years before the pandemic (only 2% of the LLMs report a difference lower than –1 standard deviation). [^26] Figure 6 in Appendix E presents the placebo maps for the single pre-treatment years (2018 and 2019). In this instance as well, very few LLMs exhibit a difference lower than –1 standard deviation between the forecasted and observed math scores prior to the pandemic (4.3% of LLMs in 2018 and 5.7% of LLMs in 2019). Furthermore, the LLMs with extreme values tend to be the smallest ones, which are also the most volatile.
Finally, Figure 3 reports a diagnostic test for the performance of the MLCM with random forest. The figure illustrates that the distribution of the placebo forecasting errors is approximately normal and centered around zero, not only for the temporal average but also for the single pre-treatment years. This demonstrates that the ATE estimator is nearly unbiased even when applied on very short panels and without relying on post-treatment information.
Figure 2: Unit-level panel placebo test—Temporal average ‘treatment’ effects. Note: the standard deviation (SD) refers to the standardized pre-COVID mean. Estimation via the MLCM with random forest.
@@ 4.3.2 Analysis Stage
Figure 4 shows the local effects of the pandemic on math scores of fifth-grade students across Italian LLMs and reports the ATE. These estimates come from the best-performing technique of the Design stage, the MLCM with random forest.
@@
The first main insight is that the COVID-19 shock engendered major learning losses in Italy. The ATE points to a decrease in the math score of –0.7608 standard deviation (SD) with respect to the pre-pandemic average, and this effect is strongly statistically significant with 95% confidence intervals (–0.7976; –0.6884). This result is in line with new micro-level studies documenting large negative impacts on educational outcomes in Italy (Battisti and Maggio,, 2023; Carlana et al.,, 2023).
As shown above, some of the smallest LLMs (the most volatile) were imprecisely estimated in the placebo exercise. Consequently, since equal weight is attributed to all treated units, the presence of these LLMs might affect the accuracy of the ATE estimate. In general, to address this potential concern, we recommend a sensitivity analysis that re-estimates the ATE following the exclusion of units with the poorest fit in the pre-treatment period. In our application, we removed the top 1%, 2%, and 5% of the most imprecisely estimated LLMs in the placebo tests. The resultant ATE estimates were -0.7500, -0.7568, and -0.7341, respectively—each closely aligning with the ATE estimate of -0.7608.272727This sensitivity test mirrors the approach proposed by Abadie et al., (2010) for the synthetic control method, which involves removing the untreated units that were imprecisely estimated from the in-space placebo test.
@@
The second key finding is that this average impact masks considerable heterogeneity across the Italian territory. Some LLMs experienced drops in students’ performances much larger than the ATE. In particular, the MLCM detects learning losses of more than one standard deviation in clusters of local economies mostly, but not exclusively, located in the South of Italy (Campania, Apulia, Molise, Calabria, Sicily, and Sardinia), and drops in math scores greater than two SD in scattered areas predominantly concentrated in Central and Southern regions.
@@
Figure 5 presents data-driven CATEs estimated with the regression tree analysis. [^28] For this analysis, we selected, on the basis of domain knowledge, a set of LLM-level variables potentially associated with the estimated treatment effects, including, among others, pre-pandemic variables such as the per capita income, the unemployment rate, the Gini index, the share of university graduates, and two proxies of access to distance learning that capture the digital divide across areas. We also included excess deaths registered during the first wave of the pandemic (Cerqua et al.,, 2021). The full list of predictors included in the regression tree CATE model can be found in Table 14 in Appendix E. The algorithm selects only three predictors to construct the tree. The analysis reveals that the most substantial learning losses (1.22 SD below the pre-COVID baseline) occurred in LLMs characterized by an unemployment rate equal to or above 10.19%, a percentage of university graduates below 26%, and a Gini index equal to or above 0.42. Conversely, the territories with an above-average treatment effect (–0.56 SD) exhibit an unemployment rate below 10.19% and a percentage of university graduates equal to or above 23%. Finally, all the CATEs reported in the tree are statistically significant.
@@
Figure 5: Data-driven CATEs. Note: The values within the terminal nodes report the CATEs (measured in SD) for all LLMs grouped within that node. Lower and upper bounds refer to the 95% confidence interval and are reported in parentheses.
@@
Given that areas with higher unemployment, greater inequality, and lower education rates at baseline have experienced disproportionate impacts, the pre-existing gaps across the country have been amplified. If such gaps persist, it is likely that these educational inequalities will, in turn, act as a catalyst for future economic inequality, leading to widened territorial disparities in the medium and long term. In summary, we can reasonably anticipate that, in the absence of counterbalancing policy efforts, the pandemic will exacerbate long-standing inequalities that predate COVID-19.
@@ 5 Conclusions
We propose a new causal panel data method based on flexible counterfactual forecasting with machine learning. The MLCM can employ any off-the-shelf ML technique and enables the estimation of policy-relevant causal parameters in evaluation settings with short and long panels and without a control group. The method is embedded within the potential outcomes framework and based on panel CV for estimation and bootstrap methods to conduct inference. To illustrate the potential of the MLCM, we presented numerical studies and an empirical analysis on the effects of the COVID-19 crisis on education in Italy, which revealed a generalized drop in students’ performance and significant inequalities in learning losses across the Italian territory. The companion R package MachineControl provides an easy-to-use implementation of the proposed approach. Large-scale shocks, international economic policies, nationwide policy changes, regional programs and local interventions engendering substantial interaction between units, are all real-word scenarios in which a suitable control group may not exist. In such cases, researchers can harness the MLCM, which complements the existing econometric toolbox for causal inference and program evaluation.
Appendix A Bootstrap inference for ATE and CATE confidence intervals
@@
Following recent work on causal machine learning estimators for panel data (Viviano and Bradic,, 2023), inference on causal effects estimated with the MLCM is performed by block-bootstrap. Since we are dealing with a panel dataset that has an inherent temporal structure, resampling the unit-time pairs independently in a context with dependent data would lead to data leakage. Therefore, we sample the units instead of the unit-time pairs so that if a unit is sampled, its entire evolution over time is sampled as well. This way, the inference is more robust to misspecification issues. This idea is in the spirit of the original block-bootstrap introduced by Carlstein, (1986) and Kunsch, (1989) that consists in dividing a time series in blocks (overlapping or non-overlapping) and then sampling the blocks with replacement. Since we have a panel dataset, we can consider each unit to form one block. The full algorithm to derive block-bootstrap confidence intervals in our setting is detailed in Algorithm 1 below.
@@
3:     Only for the resampled units, retain the tuple $(Y_{i,t}^{*},\mathcal{Y}_{i,t-1}^{(p)*}(0),\mathbf{X}_{i,t}^{*},\mathcal{X}_{i,t-1}^{(q)*})^{(b)}$ for all $t=1,\dots,T$ (pre- and post-intervention)
4:     Compute the bootstrap replication of the ATE, i.e., $\hat{\tau}^{*}_{t}(b)$ by repeating the Design stage (steps 3 to 5) and the Analysis stage (steps 6 to 9) of the MLCM outlined in Section 3
@@
6:  Let $\widehat{G}(c)=\#\{\hat{\tau}^{*}_{t}(b)\leq c\}/B$ be the empirical cumulative distribution function of the $B$ bootstrap replications. Compute a $(1-\alpha)$ percentile interval as $[\widehat{G}^{-1}(\alpha/2),\widehat{G}^{-1}(1-\alpha/2)]$.
Algorithm 1 Block-bootstrap algorithm for ATE
@@
3:     Only for the resampled units, retain the tuple $(\hat{\tau}^{*}_{t},\mathbf{X}_{i,t}^{*})_{d}^{(b)}$
4:     Compute the bootstrap replication of CATE in the terminal node $d$ by averaging the ATEs of the resampled units, i.e., $\hat{\tau}^{*}_{t}(d)$.
@@
7:  Let $\widehat{G}(c)=\#\{\hat{\tau}^{*}_{t}(d)\leq c\}/B$ be the empirical cumulative distribution function of the $d$-group CATE in the $B$ bootstrap replication. Compute a $(1-\alpha)$ percentile interval as $[\widehat{G}^{-1}(\alpha/2),\widehat{G}^{-1}(1-\alpha/2)]$.

### v2

@@ 4.3.1 Design Stage
Table 3 below reports the average performance—in terms of panel CV MSE—of the four selected ML algorithms in forecasting the standardized math test score of fifth-grade students across the pre-treatment test sets (2018 and 2019). [^26] A replication notebook of the main results reported in this section is publicly available here.
The best-performing MLCM version is the one using the random forest with twenty variables and 1/3 of them, i.e., six, as candidates at each split. Overall, the fully non-linear random forest and boosting fare better than the linear ML methods, suggesting the presence of non-linearities in the data-generating process. Moreover, the non-linear methods outperform LASSO and Partial Least Squares in all cases, and their performance is stronger in the presence of very few variables and less deteriorated by the inclusion of variables with poor predictive power (cf. Table E.2 in Supplemental Appendix E). See Table E.3 in Supplemental Appendix E for a detailed description of the 20 variables with the highest importance score attributed by the preliminary forest).
@@
Panel A – Performance
MLCM using: Panel CV-MSE
@@
Panel B – Placebo estimates with Random Forest
@@
[…]
Notes: The panel CV procedure has selected the 20 most predictive variables from the dataset for all methods. In addition, for boosting, it has selected 2,000 trees, 2 as the maximum depth of each tree, 5 as the minimum number of observations in terminal nodes and 0.005 for the learning rate. The other selected hyperparameters are: 0.3 as the penalty parameter (LASSO), 9 as the number of components (PLS), and 6 for the number of variables randomly sampled as candidates at each split (random forest). Lower and upper bounds of placebo ATE estimates refer to the 95% confidence intervals and are reported within brackets.
Panel CV on pre-treatment periods implicitly enables in-time placebo analysis along the lines of Bertrand et al. (2004) and Liu et al. (2022). In-time placebos are performed on the same pool of treated units where we “fake” that the treatment occurred at time $t_{0}-1$ and use only information up to $t_{0}-1$ to forecast the counterfactuals at time $t_{0}$. Since we know what the real values at time $t_{0}$ are, we can shift the treatment date artificially backward in time to assess the difference with the main estimates and evaluate the forecasting accuracy and unbiasedness of our estimator. As reported in the table, the placebo ATEs are indistinguishable from zero and statistically insignificant both in 2018 (ATE: 0.0554, with 95% confidence intervals (–0.1182; 0.2765) and 2019 (ATE: 0.0254, with 95% confidence intervals (–0.0273; 0.0728). [^27] The counterfactual estimates exhibit substantial similarity regardless of whether a preliminary random forest or an alternative preliminary procedure is employed, such as boosting or selecting covariates based solely on the highest absolute correlation coefficients (e.g., the correlation between the counterfactual estimates and those obtained using default hyperparameter values is 0.9985). Detailed results are available upon request.
In addition, the unit-level placebo map reported in Figure 2 depicts the temporal average (for the years 2018 and 2019) individual ‘treatment’ effects estimated with the best-performing ML algorithm, random forest, and shows that almost all LLMs exhibit no trace of significant differences between the forecasted and observed math scores in the years before the COVID-19 pandemic. In particular, only 2% of the LLMs report a difference lower than –1 standard deviation. [^28] Figure E.2 in Supplemental Appendix E presents the placebo maps for the single pre-treatment years (2018 and 2019). In these instances as well, few LLMs exhibit a difference lower than –1 standard deviation between the forecasted and observed math scores prior to the pandemic (4.3% of LLMs in 2018 and 5.7% of LLMs in 2019). Furthermore, the LLMs with extreme values tend to be the smallest ones, which are also the most volatile.
Figure 2: Unit-level panel placebo test—Temporal average ‘treatment’ effects. Notes: the standard deviation (SD) refers to the standardized pre-COVID mean. Estimation via the MLCM with random forest.
Finally, Figure 3 reports a diagnostic test for the performance of the MLCM with random forest. The figure illustrates that the distribution of the placebo forecasting errors is approximately normal and centered around zero, not only for the temporal average but also for the single pre-treatment years. This demonstrates that the ATE estimator is nearly unbiased even when applied on very short panels and without relying on the use of post-treatment information.
@@ 4.3.2 Analysis Stage
Figure 4 shows the local effects of the pandemic on math scores of fifth-grade students across Italian LLMs and reports the ATE. These estimates come from the best-performing technique of the Design Stage, the MLCM with random forest. [^29] Figure E.3 in Supplemental Appendix E provides SHAP values for this model for the counterfactual 2020 forecasts.
The first insight is that the COVID-19 shock engendered major learning losses in Italy. The ATE points to a decrease in the math score of –0.7608 standard deviation (SD) with respect to the pre-pandemic average, and this effect is strongly statistically significant (95% confidence intervals: –0.7976; –0.6884). This result is in line with micro-level studies documenting large negative impacts on educational outcomes in Italy (Battisti and Maggio, 2023; Carlana et al., 2023).
As shown above, some of the smallest LLMs (the most volatile) were imprecisely estimated in the placebo exercise. Since equal weight is attributed to all treated units, the presence of these LLMs might affect the accuracy of the ATE estimate. To address this potential concern, we recommend a sensitivity analysis that re-estimates the ATE following the exclusion of units with the poorest fit in the pre-treatment period. Here, we removed the top 1%, 2%, and 5% of the most imprecisely estimated LLMs in the placebo tests. The resultant ATE estimates were -0.7500, -0.7568, and -0.7341, respectively—each closely aligning with the ATE estimate of -0.7608.
@@
The second key finding is that this average impact masks considerable heterogeneity across the Italian territory. Some LLMs experienced much larger drops in students’ performances. In particular, the MLCM detects learning losses of more than one standard deviation in clusters of local economies mostly located in the South of Italy (Campania, Apulia, Molise, Calabria, Sicily, and Sardinia), and drops in math scores greater than two SD in scattered areas predominantly concentrated in Central and Southern regions.
@@
Figure 5 presents data-driven CATEs estimated with the regression tree analysis. [^30] For this analysis, we selected a set of LLM-level variables including, among others, pre-pandemic variables such as the per capita income, the unemployment rate, the Gini index, the share of university graduates, two proxies of access to distance learning that capture the digital divide across areas, and excess deaths registered during the first wave of the pandemic (Cerqua et al., 2021). The full list can be found in Table LABEL:tabapp:cate_vars in Supplemental Appendix E. The algorithm selects only three predictors to construct the tree. The analysis reveals that the most substantial learning losses (1.22 SD below the pre-COVID baseline) occurred in LLMs characterized by an unemployment rate equal to or above 10.19%, a percentage of university graduates below 26%, and a Gini index equal to or above 0.42. Conversely, the territories with an above-average treatment effect (–0.56 SD) exhibit an unemployment rate below 10.19% and a percentage of university graduates equal to or above 23%. Finally, all the CATEs reported in the tree are statistically significant.
@@
Figure 5: Data-driven CATEs. Notes: The values within the terminal nodes report the CATEs (measured in SD) for all LLMs grouped within that node. Lower and upper bounds refer to the 95% confidence interval and are reported in parentheses.
@@
Since areas with higher unemployment, greater inequality, and lower education rates at baseline have experienced disproportionate impacts, the pre-existing gaps across the country have been amplified. If such gaps persist, it is likely that they will act as a catalyst for future economic inequality, leading to widened territorial disparities in the medium and long term. In summary, we can anticipate that, in the absence of counterbalancing policy efforts, the pandemic will exacerbate long-standing inequalities that predate COVID-19.
@@ 5 Conclusions
Identifying causal effects is challenging without a credible control group. We overcome this challenge by proposing a new method based on counterfactual forecasting with machine learning. The MLCM can employ any off-the-shelf ML algorithm to estimate policy-relevant causal parameters in a wide variety of panel settings, including very short panels, without relying on untreated units. To illustrate its potential, we presented numerical studies, a replication of the minimum wage application in Callaway and Sant’Anna (2021) without using untreated units, and an empirical analysis on the effects of the COVID-19 crisis on education in Italy. The companion R package MachineControl provides an easy-to-use implementation of the proposed approach. Large shocks, international economic policies, nationwide policy changes, and regional programs engendering substantial interaction between units, are all real-world scenarios where a control group may not exist. In such cases, researchers can harness the MLCM, which complements the existing econometric toolbox for causal inference and policy evaluation.
Appendix A Causal estimands in settings with interference between units
Compared to Case (i), in the scenario with violations of the no-interference assumption, the causal estimands are different. In this section, we define the main causal estimands for the setup in which only a subgroup of units is treated, but there are spillover and general equilibrium effects that prevent using the set of untreated units to build a valid control group (Case (ii) defined above). We can first define the ATT across the treated units at a given point in time.
Definition 9.
Definition 10.
Appendix B Identification proof
We show here the general identification proof for the causal estimands defined in Section 2.2.
Proof.
As in the main text, the proof is organized by distinguishing between $k=1$ and $k=2$.
  • $k=1$. Let $\operatorname{Y}_{i,t}(0)$ follow Equation (2). In the expression $\tau_{i,t_{0}+1}=\operatorname{Y}_{i,t_{0}+1}(1)-\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$, the first term $\operatorname{Y}_{i,t_{0}+1}(1)$ is the observed outcome under the policy and thus is immediately identified. Since $Q=\max{\{0,q\}}=q$, the second term can be written as
    $\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+1}(0)|\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+1}^{(q)}]$ Def. of $\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$
    $=\operatorname{E}[h(\operatorname{Y}_{i,t_{0}}^{(p)}(0),\mathbf{X}_{i,t_{0}+1}^{(q)})|\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+1}^{(q)}]$ A.4 - Eq.(2)
    $=h(y_{i,t_{0}}^{(p)}(0),\mathbf{x}_{i,t_{0}+1}^{(q)})=y_{i,t_{0}+1|t_{0}}(0).$
    The last expression follows from the fact that $h(\cdot)$ is deterministic once we observe $\operatorname{Y}_{i,t_{0}}^{(p)}(0)=y_{i,t_{0}}^{(p)}(0)$ and $\mathbf{X}^{(q)}_{i,t_{0}+1}=\mathbf{x}^{(q)}_{i,t_{0}+1}$. As a result, $\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$ is identified from observed data, and so is $\tau_{i,t_{0}+1}$. Notice that in this proof, we never used the linearity of the potential outcome model. Thus, the proof at $k=1$ is the same even when $h(\cdot)$ is non-linear.
  • $k=2$. In the expression $\tau_{i,t_{0}+2}=\operatorname{Y}_{i,t_{0}+2}(1)-\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$, the first term $\operatorname{Y}_{i,t_{0}+2}(1)$ is the observed outcome under the policy. Thus, the identification of the causal effect depends solely on the second term, $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$. Since $Q=\max{\{1,q\}}$, we now have two possible cases: either $q\geq 1$, simplifying to $Q=q$ or $q=0$, in which case $Q=1>q$. Assuming the latter, we now distinguish between a linear and a non-linear model specification:
    Linear case. Assume the following linear model $h(\operatorname{Y}_{i,t-1}^{(p)}(0),\mathbf{X}_{t}^{(q)})=\sum_{j=0}^{p}b_{1}^{(j)}\operatorname{Y}_{i,t-1-j}(0)+\sum_{j=0}^{q}\mathbf{b_{2}}\mathbf{X}_{i,t-q}$. The term $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ can then be written as,
    $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+2}(0)|\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]\hskip 143.0pt\text{Def. $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$}$
[…]
    $=\operatorname{E}[b_{1}^{(0)}\operatorname{Y}_{i,t_{0}+1}(0)+\sum_{j=1}^{p}b_{1}^{(j)}\operatorname{Y}_{i,t_{0}+1-j}+\mathbf{b_{2}}\mathbf{X}_{i,t_{0}+2}|\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]\hskip 15.0pt\text{Linearity }$
[…]
    $=b_{1}^{(0)}y_{i,t_{0}+1|t_{0}}(0)+\sum_{j=1}^{p}b_{1}^{(j)}y_{i,t_{0}+1-j}+\mathbf{b_{2}}\mathbf{x}_{i,t_{0}+2}.$
    Non-linear case. The identification of $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ can be based on the formulation for the post-intervention potential outcome model defined by Equation (3),
    $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+2}(0)|\operatorname{Y}_{i,t_{0}}(0),\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]\hskip 105.0pt\text{Def. $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$}$
[…]
    $=\operatorname{E}[g_{2}(h(\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+1}),\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+2})|\operatorname{Y}_{i,t_{0}}^{(p)},\mathbf{X}_{i,t_{0}+2},\mathbf{X}_{i,t_{0}+1}]$
    $=g_{2}(y_{i,t_{0}+1|t_{0}}(0),y_{i,t_{0}}^{(p)},\mathbf{x}_{i,t_{0}+2}).$
    The last expression follows from the fact that $g_{2}(\cdot)$ is deterministic once we observe $\operatorname{Y}_{i,t_{0}}=y_{i,t_{0}}$, $\mathbf{X}_{i,t_{0}+1}=\mathbf{x}_{i,t_{0}+1}$ and $\mathbf{X}_{i,t_{0}+2}=\mathbf{x}_{i,t_{0}+2}$. In addition, even though Assumption 3 was not explicitly mentioned in this part of the proof, notice that it is implicit in Assumption 5, as by Equation (3), the potential outcomes never depend on future covariates.
The proof for a generic time $t_{0}+k$ follows analogously by induction. ∎
Appendix C Bootstrap inference for ATE and CATE confidence intervals
@@
Following recent work on causal machine learning estimators for panel data (Viviano and Bradic, 2023), inference on causal effects estimated with the MLCM is performed by block-bootstrap. Since we are dealing with a panel dataset that has an inherent temporal structure, resampling the unit-time pairs independently in a context with dependent data would lead to data leakage. Therefore, we sample the units instead of the unit-time pairs so that if a unit is sampled, its entire evolution over time is sampled as well. This way, the inference is more robust to misspecification issues. This idea is in the spirit of the original block-bootstrap introduced by Carlstein (1986) and Kunsch (1989) that consists in dividing a time series in blocks (overlapping or non-overlapping) and then sampling the blocks with replacement. Since we have a panel dataset, we can consider each unit to form one block. The full algorithm to derive block-bootstrap confidence intervals in our setting is detailed in Algorithm 3 below.
@@
3:     Only for the resampled units, retain the tuple $({\operatorname{Y}}_{i,t}^{(p)}(0),\mathbf{X}_{i,t}^{(q)})^{(b)}$ for all $t=1,\dots,T$ (pre- and post-intervention)
4:     Compute the bootstrap replication of the ATE, i.e., $\hat{\tau}^{(b)}_{t}$ by repeating the Design stage (steps 3 to 5) and the Analysis stage (steps 6 to 9) of the MLCM outlined in Section 3
@@
6:  Let $\widehat{G}(c)=\#\{\hat{\tau}^{(b)}_{t}\leq c\}/B$ be the empirical cumulative distribution function of the $B$ bootstrap replications. Compute a $(1-\alpha)$ percentile interval as $[\widehat{G}^{-1}(\alpha/2),\widehat{G}^{-1}(1-\alpha/2)]$.
Algorithm 3 Block-bootstrap algorithm for the ATE
@@
3:     Only for the resampled units, retain the tuple $(\hat{\tau}_{i,t},\mathbf{X}_{i,t})_{d}^{(b)}$
4:     Compute the bootstrap replication of CATE in the terminal node $d$ by averaging the ATEs of the resampled units, i.e., $\hat{\tau}_{t}(d)$.
@@
7:  Let $\widehat{G}(c)=\#\{\hat{\tau}_{t}(d)\leq c\}/B$ be the empirical cumulative distribution function of the $d$-group CATE in the $B$ bootstrap replication. Compute a $(1-\alpha)$ percentile interval as $[\widehat{G}^{-1}(\alpha/2),\widehat{G}^{-1}(1-\alpha/2)]$.